Express the following linear equations in the form ax + by + c = 0 and indicate the values of a, b and c in each case:

  y – 2 = 0

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


y – 2 = 0


0x + 1y + (–2) = 0

Now comparing 0x + 1y + (–2) = 0with ax + by + c = 0

We get, a = 0

b = 1

c = –2

Thanks